OpenStudy (anonymous):

The derivative of h(x) is given by the function h '(x) = x + 2cosx. Use your calculator to determine the number of times the graph of h(x) changes concavity on the interval [0, ]. Justify your answer using a complete sentence.

OpenStudy (anonymous):

[0,4PI]

OpenStudy (anonymous):

derive it two times and do it equal zero

OpenStudy (anonymous):

= 1 - 2sin(X)

OpenStudy (anonymous):

dont get the second part

OpenStudy (anonymous):

w8 nvm is it 2?

OpenStudy (anonymous):

concavity is tested by y'' test

OpenStudy (anonymous):

0=1-2sin(x) -1=-2sin(x) 1/2=sin(x)

OpenStudy (anonymous):

happens at pi/6

OpenStudy (anonymous):

only once?

OpenStudy (anonymous):

nope it says in the interval of 0 to 4pi

OpenStudy (anonymous):

so two revolutions

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so it's pi/6, 5pi/6, 13pi/6, 17pi/6

OpenStudy (anonymous):

using a calculator if you look at the graph

OpenStudy (anonymous):

k thanks

OpenStudy (anonymous):

what do you mean by 2 revolutions?

OpenStudy (anonymous):

you'd have to integrate up one to actually see the graph of the function but the inflection points http://www.wolframalpha.com/input/?i=x%5E2%2F2%2B+2*sinx+from+0+to+4pi you can see at certain points it switches from concave up to concave down

OpenStudy (anonymous):

2 revolutions around the unit circle

OpenStudy (anonymous):

4pi =720 degrees

OpenStudy (anonymous):

alright

OpenStudy (anonymous):

thanks again

OpenStudy (anonymous):

no problm